If each of the equal side of an isosceles triangle is 10, and the base is 16, what is the area of the triangle?
(a) 39 (b) 48 (c) 80 (d) 96 (e) 160

Answer by rfer(16322)

B) 48
10^2-8^2=36
sq rt 36=6
(6*16)/2=48